# Implementation Study of Cost-Effective Verification for Pietrzak’s Verifiable Delay Function in Ethereum Smart Contracts

### v3

@@ Abstract
Verifiable Delay Function (VDF) is a cryptographic concept that ensures a minimum delay before output through sequential processing, which is resistant to parallel computing. Among the two well-known VDF protocols, Wesolowski and Pietrzak VDF, we focus on the Pietrzak VDF due to its computational efficiency and suitability for blockchain environments. Pietrzak’s approach uses a recursive proof verification with the halving protocol, offering a practical alternative despite the longer proof length than Wesolowski’s approach. Given the scarcity of research on practical VDF verification implementation, especially within smart contracts, this paper aims to implement cost-effective verification for the Pietrzak VDF in an Ethereum-based environment without compromising the VDF verification’s integrity and reliability. Firstly, we propose generalized proof generation and verification algorithms for potential efficiency improvement. Secondly, we categorize and measure the gas cost of each part in a transaction for VDF verification. Thirdly, based on the analysis, we theoretically predict the optimized proof construction. Finally, we demonstrate the theoretical prediction matches the implementation results. Furthermore, our research shows that the proof length of the Pietrzak VDF is generated under 8 KB with the 2048-bit RSA key length, much smaller than the previous expectation. This implies that the Pietrzak VDF can be practically used for cryptographic applications on blockchains.
@@ 1 Introduction
footnotetext: We express our special thanks to Prof. Seungjoo Kim at Korea University for his invaluable advice on cryptographic algorithms.
In blockchain technology, one critical element is the generation and use of cryptographically secure random numbers. These secure random numbers play a pivotal role in various applications of blockchain; for example, in consensus, random numbers can ensure fair selections of sequencers and block proposers based on a probabilistic process. Random numbers are essential for ensuring anonymity and serve as the basis for cryptographic techniques. Despite the significance of secure random numbers, incorporating them into smart contracts results in challenges. Due to the lack of random number methods in the smart contract interface, individuals previously used to apply future block data (i.e. the blockhash of a future block) as an ad-hoc random number. It can be manipulated by a block generator. Hence, leveraging ad-hoc random numbers is not a suitable approach for the present situation.
In this context, two notable techniques have emerged: Verifiable Random Functions (VRFs) and Verifiable Delay Functions (VDFs). VRFs are cryptographic primitives that produce a random number along with proof of its correctness, allowing anyone to verify the number’s authenticity without compromising its unpredictability. Several commercial random number generating services using VRF, such as Chainlink VRF, are already available. As part of their operation, a node acting as an oracle transmits a randomly generated number from off-chain nodes to a blockchain.
On the other hand, VDFs are designed to produce a unique output that requires a predetermined amount of time to compute, but whose result can be efficiently verified by others. Therefore, VDFs do not directly generate a random number. Rather, VDFs support random number generation of the commit-reveal scheme by utilizing their time-lock property. The most critical weakness of the commit-reveal scheme is that all the participants must reveal to generate a secure random number. VDFs’ time-lock property ensures that the random outcome cannot be predicted or accelerated by computational power in case of any absence. It makes VDFs practical for applications that benefit from a guaranteed delay in randomness generation. Financial penalties alone in the commit-reveal scheme may not ensure the reliable generation of random numbers. Integrating VDFs with it, restructured as a commit-reveal-recover mechanism, can significantly enhance security.
Figure 1: In our model, a user sends a VDF proof after the Pietrzak VDF evaluation. The primary issue is the significant gas usage for the VDF verification. The secondary issue is the length of the VDF proof as the Pietrzak VDF’s length increases as the time delay increases.
Pivotal advancements in VDFs have been made through the research of Pietrzak [1] and Wesolowski [2], who each proposed practical implementations of VDFs. Despite these theoretical developments, there remains a notable gap in research regarding the application of VDFs within blockchain environments. The challenge lies not only in the theoretical construction of VDFs, but also in adapting these constructions to the resource usage context of blockchain systems. For example, the Pietrzak VDF uses the halving protocol that makes the length of proof proportional to the delay time. It makes utilizing the Pietrzak VDF considered unrealistic in blockchain [3]. The Wesolowski VDF needs a hash-to-prime oracle that is not supported in most blockchains.
When implementing VDFs, the process can broadly be divided into three parts: evaluation, proof generation, and verification. In the context of blockchain, the verification component must be implemented on-chain. This segmentation is due to our belief that in typical blockchain applications, both evaluation and proof generation processes are performed off-chain, mainly due to their computational intensity and the associated costs when executed on a blockchain network.
Our goal has been to efficiently implement VDF verification within smart contracts as Figure 1’s illustration. Specifically, we have targeted the Pietrzak VDF model. There are two reasons: Firstly, the Wesolowski VDF requires a hash-to-prime function that is challenging to implement on smart contracts. Secondly, we conjectured that optimization can mitigate the long-proof length issue partially. Based on this algorithm choice, consequently, this research demonstrated that the Pietrzak VDF verification can be utilized in the EVM-compatible smart contract environment through various cost-saving analysis and experiments.
@@
  • We proposed generalized algorithms for proof generation and verification of the Pietrzak VDF for potential efficiency enhancement.
@@
The rest of the paper is organized as follows: Section 2 introduces related studies about VDFs and their implementation. Section 3 describes the proposed target algorithms and our implementation environment. Section 4 illustrates the data structures of proofs and analyzes gas costs for the data and its dispatching cost. Section 5 studies arithmetic operations in the verification process and their gas costs. In Section 6, based on the previous sections’ outcomes, we present theoretical analysis and the cost-effective implementation. Section 7 discusses this research’s contributions, limitations, and future directions. Finally, Section 8 concludes the paper.
2 Background
This section introduces VDFs, cryptographic algorithms crucial for blockchain applications and secure timestamping. It also reviews related works that inform the VDF implementations discussed in this paper.
2.1 Verifiable Delay Functions
VDFs are cryptographic protocols crafted to ensure that their execution takes a predefined minimum amount of time, regardless of the computational power applied. Here we introduce the core mechanisms that define VDFs and the core properties that VDFs get from the elements.
2.1.1 Definition and Fundamental Properties
A VDF consists of three key algorithms that facilitate one round: setup, evaluation, and verification.
  • Setup $(\text{Setup}(1^{\lambda})\rightarrow(\mathbb{G},T))$: Initializes the VDF with public parameters. Here, $1^{\lambda}$ represents the security parameter indicating the required security level, $\mathbb{G}$ is a cryptographic group used in computations, and $T$ is the time delay parameter that sets the duration of the computation.
  • Evaluation $(\text{Evaluate}(\mathbb{G},T,X)\rightarrow(Y,\pi))$: For a given input $X$, the VDF computes an output $Y$ over time $T$, which is deliberately designed to be sequential, preventing acceleration through parallel processing. The output includes a proof $\pi$ that validates the computation.
  • Verification $(\text{Verify}(\mathbb{G},T,X,Y,\pi)\rightarrow\{\textsf{true},\textsf{false}\})$: Enables any observer to quickly and efficiently verify the correctness of the output $Y$ and its proof $\pi$ without undergoing the time-intensive computation themselves. A return of true confirms the validity of the output, while false indicates an error.
From the above algorithms, VDFs must get the three primary attributes [4]:
  • Sequentiality in evaluation: The computation must take a fixed minimum time, irrespective of the available computational resources.
  • Efficiency in verification: Although the computation itself is intentionally delayed, the verification process is designed to be quick and require minimal resources.
  • Uniqueness in output: Each input to a VDF yields a unique, unpredictable output, ensuring reliability and consistency in applications.
2.2 Feasibility of On-Chain VDF Verification
Both Wesolowski [2] and Pietrzak [1] VDFs operate by transforming a value $x$ into $y=x^{2^{T}}\mod N$, where $N$ is an RSA modulus, although their verification methods differ. Scalability and implementation costs are significant challenges in blockchain systems. Employing the Fiat-Shamir heuristic, both methods facilitate non-interactive verification, suitable for blockchain environments. However, the computational intensity and memory demands for managing large numeric values pose substantial hurdles.
Wesolowski’s VDF maintains a constant proof length ($O(1)$), independent of the time delay parameter $T$, making it particularly advantageous where proof size is a concern. The implementation on blockchain platforms like Ethereum is particularly challenging due to the high cost associated with verifying or finding primes for large numbers, such as 2048 bits. This necessity significantly increases the implementation cost, making full on-chain execution unfeasible.
Pietrzak’s VDF does not require operations like hash-to-prime, simplifying its on-chain implementation. The protocol involves a halving process, where $T$ is divided by two until it becomes one, reducing the complexity of computations with each iteration. However, the proof length and cost are influenced by the time delay variable $T$, and the proof length grows with $O(\log_{2}(T))$. This characteristic can result in verifications that exceed the size limit for feasible blockchain implementation. The research anticipated that the length of the proof will approximate 40 KB [3].
2.3 Related Works
In this subsection, we review the critical advancements in VDFs, from their theoretical foundations to practical implementations related to this study.
Theoretical research on VDFs. The concept of VDFs has garnered significant attention for its potential to enable new applications in cryptography and blockchain technology. The foundational work by Rivest, Shamir, and Wagner on "Time-Lock Puzzles and Timed-Release Crypto" [5] laid the groundwork for the development of protocols that secure data until a certain amount of time has passed. This seminal paper introduced the concept of time-lock puzzles, a precursor to VDFs, demonstrating how cryptographic operations could be used to create a delay mechanism that is practically infeasible to bypass without solving the puzzle.
Building upon these principles, Pietrzak [1] and Wesolowski [2] independently proposed two distinct approaches to VDFs, each with unique advantages in terms of efficiency and security. Pietrzak’s method focuses on using iterated sequential squaring in a group of unknown order, providing a balance between computation time and verification efficiency. Wesolowski’s approach, on the other hand, introduces an elegant solution that reduces the verification complexity, making it more practical for applications that require fast and efficient verification processes.
Further advancing the field, Dan Boneh et al. [6] have conducted extensive research on optimizing VDFs, addressing critical aspects such as the construction of more efficient protocols and the reduction of computational requirements for both generation and verification phases. Their work not only highlights the theoretical advancements in VDF technology but also emphasizes the practical implications for secure, decentralized systems.
@@ 2.3 Related Works
Implementation of VDFs. The practical realization of VDFs has become an area of intense research and development, seeking to bridge the gap between theoretical robustness and real-world applicability. A significant contribution to this endeavor is the study by Attias et al. [3], which provides a comprehensive implementation analysis of two leading VDFs, Pietrzak and Wesolowski. Their work meticulously evaluates the performance of VDFs’ evaluation, proof generation, and verification. Their work provided us with insight into the research and optimization techniques for VDFs. However, they expected the Pietrzak VDF would not be viable for blockchains. This is contrary to our results that presented the feasibility of the Pietrzak VDF on Ethereum smart contracts.
Choi et al. [7] proposed three practical commit-reveal-recover techniques. Their protocols are crucial for establishing the commit-reveal-recover scheme, which mostly existed at the conceptual level. They incorporated the Wesolowski VDF into Ethereum smart contracts for the purpose of demonstration. Our research covers VDF verification, which corresponds to the Proof-of-Exponentiation (PoE) stated in their work. Their PoE gas usage is approximately 2.3 million, which is comparable to the gas consumption of our work with the time delay variable of $2^{22}$. Due to the Wesolowski VDF not having increased calculation requirements with greater time delays, our approach becomes less efficient when the time delay variable exceeds $2^{23}$. Nevertheless, the Wesolowski VDF encounters an issue when it comes to fully implementing the hash-to-prime function as previously indicated.
@@ 3 Implementation Target and Environment
In this section, we provide the target algorithm and environment. Firstly, we define the Pietrzak VDF algorithms we aim to implement in this research. Secondly, we describe the experiment environment in the following sections.
3.1 Pietrzak VDF Algorithms
Our research has a model with a smart contract that supports a function that performs Pietrzak VDF verification. Referring to the Pietrzak’s original paper [1], we define the evaluation as described in Algorithm 1 and the halving protocol for proof generation in Algorithm 2.
3.1.1 VDF Evaluation
$y\equiv x^{2^{T}}\equiv x^{2^{2^{\tau}}}\mod{N}$ (1)
The VDF evaluation process as defined in Algorithm 1 is pivotal in generating a VDF output from given inputs. It basically computes Equation 1. The algorithm takes an initial value $x$, a time parameter $T$ (also notated as $2^{\tau}$), and a modulus $N$ to produce the output $y$. $N$ that is a $\lambda$-bit number is produced by an RSA setup ($N=p\cdot q$ where $p,q$ are two big prime numbers). The process involves exponentiating $x$ to $2^{T}\mod N$ to ensure the difficulty and time consumption of the evaluation are as intended. This evaluation serves as the foundation for verifying that the output can only be computed with a predefined amount of real-time computational work, which is crucial in applications requiring guarantees against parallelization.
Algorithm 1 Evaluation of the Pietrzak VDF
1:input: $x,T,N$
2:output: $y$
3:$y\leftarrow x$
4:for $k\leftarrow 1$ to $T$ do
5:     $y\leftarrow y^{2}\mod N$
6:end for
7:return $y$
3.1.2 Halving Protocol
The halving protocol detailed in Algorithm 2 optimizes the proof generation process for VDFs by reducing the verification time incrementally. At each step, the protocol halves the time parameter $T$, recalculates intermediate values $x_{i}$ and $y_{i}$, and updates these values based on a cryptographic hash function to ensure integrity. As we implement it on Ethereum, we use Keccak256 as the cryptographic hash function. This iterative reduction continues until the desired granularity is reached, ensuring that the verifier’s workload is minimized while maintaining cryptographic security.
Algorithm 2 Halving Protocol of the Pietrzak VDF
1:input: $N,x_{i},T,y_{i}$
2:output: $N,x_{i+1},T/{2^{i}},y_{i+1},\mu_{i+1}$
3:$\mu_{i}\leftarrow x_{i}^{2^{T/{2^{i}}}}$
4:$r_{i}\leftarrow\text{Hash}(x_{i}+y_{i}+v_{i})$
5:$x_{i+1}\leftarrow x_{i}^{r_{i}}\cdot v_{i}\mod N$
6:if $T/2^{i-1}$ is odd then
7:     $y_{i}\leftarrow y_{i}^{2}\mod N$
8:end if
9:$y_{i+1}\leftarrow v_{i}^{r_{i}}\cdot y_{i}\mod N$
10:return $N,x_{i+1},T/{2^{i}},y_{i+1},\mu_{i+1}$
3.1.3 Introduction of Generalized Proof Generation/Verification and Its Rationale
In the original paper [1] and the related works [6, 3], the conventional Pietrzak halving protocol is repeated until the time parameter $T$ gets 1 for proof generation. It helps the verifier consume the nearly least resources to verify VDF proofs in general systems. However, on Ethereum, the cost of resource usage, called gas, significantly differs from general environments. Ethereum Yellowpaper [8] provides predefined gas prices for transactions and opcodes executions. Especially in this context, repeating the halving protocol till $T=1$ does not guarantee the most efficient verification. Therefore, we generalized the Pietrzak verification algorithm with flexible repetition of the halving protocol. In Algorithm 4, we introduced a new parameter $\delta$. The halving protocol is repeated till the time variable becomes equal to $\delta$. Then, the verifier checks $y\stackrel{{\scriptstyle?}}{{=}}x^{2^{2^{\delta}}}$.
3.1.4 Generalized Proof Generation
In Algorithm 3, the generalized proof generation is used to construct the cryptographic proofs required for VDF verification under flexible conditions. The proof generation starts with the evaluation result and halves the exponentiation of a proof step-by-step until reaching the specified parameter $\delta$. Each halving can be tracked as intermediate values $v_{i}$ are stored as part of the proof. This mechanism allows for flexible adjustment of the verification process to accommodate different system capacities and security requirements.
Algorithm 3 Generalized Proof Generation of the Pietrzak VDF with Evaluation
1:input: $x,T,\delta,N$
2:output: $\{\pi_{i}\}_{i=1}^{\tau-\delta}$
@@ 3.1.4 Generalized Proof Generation
4:$y\leftarrow x$
5:for $k\leftarrow 1$ to $T$ do
6:     $y\leftarrow y^{2}\mod N$
7:end for
8:$(x_{1},y_{1})\leftarrow(x,y)$
9:for $i\leftarrow 1$ to $\tau-\delta$ do
10:     $v_{i}\leftarrow x^{2^{T/2^{i}}}$
11:     $r_{i}\leftarrow keccak256(x_{i}+y_{i}+v_{i})$
12:     $x_{i+1}\leftarrow x_{i}^{r_{i}}\cdot v_{i}\mod N$
13:     if $T/2^{i-1}$ is odd then
14:         $y_{i}\leftarrow y_{i}^{2}\mod N$
15:     end if
16:     $y_{i+1}\leftarrow v_{i}^{r_{i}}\cdot y_{i}\mod N$
17:     $\pi_{i}\leftarrow v_{i}$
18:end for
19:return $\{\pi_{i}\}_{i=1}^{\tau-\delta}$
3.1.5 Generalized Verification
Algorithm 4 describes the generalized verification process for VDFs on Ethereum. It utilizes the proofs generated by the generalized proof generation method to validate the sequence of operations leading to the final result. The verifier iteratively computes and checks each step of the proof against the corresponding intermediate values and the final condition $y_{\tau-\delta+1}=x_{\tau-\delta+1}^{2^{2^{\delta}}}$. This process ensures that the computation was performed correctly and meets the specified delay constraints, which is crucial for applications relying on the integrity of timed cryptographic operations.
Algorithm 4 Generalized Verification of the Pietrzak VDF
@@ 3.1.5 Generalized Verification
According to Algorithm 3, and 4, the data should be submitted are the root ($x$), the VDF evaluation result ($y=x^{2^{T}}$), the time delay variable ($T$), the VDF proof ($\{\pi_{i}\}_{i=1}^{\tau-\delta}$), and the group description ($N$) as described in Equation 2.
$\Big{(}N,x,y,\tau,\overbrace{(\pi_{1},\pi_{2},\pi_{3},...,\pi_{k})}^{\overbrace{\lfloor\log_{2}(T)\rfloor}^{\tau}-\delta}\Big{)}$ (2)
3.2 Target Time Delay and the Bit Length
Recent advances in VDF implementation using C++, as studied by Attias et al. [3], demonstrate that VDF evaluation times can range from $2^{20}$ to $2^{25}$ cycles. These findings suggest that the implementations are not only effective but also efficient in experimental setups. In the context of Ethereum, where block generation typically requires around 12 seconds and even less time in other blockchain systems, we recommend that the time delay for generating VDF-based random numbers should not exceed one minute. This recommendation aims to prevent excessive resource consumption without compromising practical utility.
The experimental results from Attias et al. [3] show that a performance duration of approximately one minute at the 2048-bit RSA key level is feasible. Consequently, we advocate for a similar time range in VDF applications. The security of our target Pietrzak VDF configuration depends on the difficulty of the prime factoring problem, akin to that in the RSA encryption system. Currently, a 2048-bit key length is deemed sufficient for maintaining security standards. However, considering potential advancements in computational capabilities, a 3072-bit key length might be necessary after ten years. Therefore, our experiments included tests with both 2048-bit and 3072-bit settings for VDFs.

### v4

# Implementation Study of Cost-Effective Verification for Pietrzak’s VDF in Ethereum Smart Contracts

@@ Abstract
Verifiable Delay Function (VDF) is a cryptographic concept that ensures a minimum delay before output through sequential processing, which is resistant to parallel computing. One of the significant VDF protocols academically reviewed is the VDF protocol proposed by Pietrzak. However, for the blockchain environment, the Pietrzak VDF has drawbacks including long poof size and recursive protocol computation. In this paper, we present an implementation study of Pietrzak VDF verification on Ethereum Virtual Machine (EVM). We found that the discussion in the Pietrzak’s original paper can help a clear optimization in EVM where the costs of computation are predefined as the specific amounts of gas. In our results, the cost of VDF verification can be reduced from 4M to 2M gas, and the proof length can be generated under 8 KB with the 2048-bit RSA key length, which is much smaller than the previous expectation.
@@ 1 Introduction
The Verifiable Delay Function (VDF)’s major role is to guarantee a specific delay for evaluation. It can play an important role in security applications. The most notable application of blockchain using VDF is a distributed randomness beacon. VDFs do not directly generate a random number. Rather, they help to guarantee the liveness of a random number generation scheme, Commit-Reveal, by recovering a random number when the reveal phase fails. This has led Ethereum, one of the largest cryptocurrencies, to include VDFs in their future roadmap for the RANDAO block proposer selection mechanism.
In recent years, several VDF mechanisms have been proposed [1, 2, 3, 4]; however, the concept of VDFs remains relatively new. This indicates that we need comprehensive research on VDFs across various dimensions. For example, recently found parallel computation mechanisms for VDF evaluation [5, 6] show that we need more research on the squaring computation. On the other hand, there is a scarcity of implementation research [7] on VDFs. In this paper, we focus on VDF implementation on EVM, one of the leading blockchain environments.
Our goal is to efficiently implement VDF verification within smart contracts, providing a foundation for EVM security applications utilizing VDFs. Specifically, we focus on the Pietrzak VDF [2] for two key reasons: First, the Wesolowski VDF [1], another major VDF, requires a hash-to-prime function, which is challenging to implement directly in smart contracts. Second, we hypothesize that the performance optimization technique discussed by Pietrzak [2] could significantly reduce the overall cost within the EVM environment. As a result of this focus, our research demonstrates through theoretical analysis and experiments that Pietrzak VDF verification can be effectively utilized within EVM-compatible smart contracts.
@@
The rest of the paper is organized as follows: Section 2 reviews related works on VDFs. Section 3 describes the target algorithm and experiment environment. In Section 4, we measure the gas costs for VDF verification on the EVM and formulate them into a function using regression. Section 5 analyzes the optimal parameters for VDF verification costs and presents experimental results that demonstrate the theoretical analysis. Finally, Section 6 concludes the paper with a discussion of further research directions. The source code for our implementation is available in the Appendix A.
2 Related Works
The concept of recursive computation for delayed output was firstly proposed as the "Time-Lock Puzzle" by Rivest et al. [8]. The Time-Lock Puzzle does not include efficient public verifiability of the solution. This was later addressed by Mahmoody et al. in their Proof of Sequential Work (PoSW) [9]. Lenstra and Wesolowski proposed Sloth, which is considered the first Verifiable Delay Function (VDF) style mechanism [3]. Boneh et al. [10] formalized the concept of VDF. Following the introduction of VDFs, various mechanisms were proposed, among which the VDFs by Wesolowski [1] and Pietrzak [2] had unique advantages in terms of efficiency and security. For the use in blockchain, Veedo and Minroot [4] were proposed for Starknet and RANDAO respectively. However, recent findings by Biryukov et al. [5] and Leurent [6] demonstrate that distributed algorithms for modular squaring exist. This implies that the fundamental assumption of using sequential squaring puzzles is not valid, highlighting the need for further research on this issue.
@@ 2 Related Works
Not only theoretical research, but also implementation studies are crucial for approaching practical applications. However, such studies appear to have been conducted in a rather limited scope. A significant contribution is the study by Attias et al. [7], which provides a comprehensive implementation analysis of two leading VDFs, Pietrzak and Wesolowski. Their work meticulously evaluates the performance of VDFs’ evaluation, proof generation, and verification. Their work provided us with insight into the research and optimization techniques for VDFs. However, they expected the Pietrzak VDF would not be viable for blockchains. This is contrary to our results that presented the feasibility of the Pietrzak VDF on Ethereum smart contracts. Choi et al. [11] proposed three practical commit-reveal-recover protocols. They incorporated the Wesolowski VDF into Ethereum smart contracts for the purpose of demonstration. Our research covers VDF verification, which corresponds to the Proof-of-Exponentiation (PoE) stated in their work.
@@ 3 Implementation Target and Environment
In this section, we provide the target algorithm and environment. Firstly, we define the Pietrzak VDF verification algorithm we aim to implement. Secondly, we describe the experiment environment for the cost measurement.
3.1 Pietrzak VDF verification
Algorithm 1 outlines the Pietrzak VDF verification process. Given that a prover insists the VDF evaluation $x^{2^{T}}=y$ with the time delay parameter $T=2^{\tau}$, the set of halving proof $\{\pi_{i}\}_{i=1}^{\tau}$, and the modulus $N$. It repeats the halving protocol (lines 5 – 13) until the verifier can simply check the result by squaring the base. The halving protocol reduces the number of exponentiations required for verification by half. The halving protocol outputs $x_{1},y_{1}$ and the proof $\pi$ reasoning $x_{1}^{2^{2^{\tau-1}}}=y_{1}$. Therefore, the verifier essentially needs to repeat the halving protocol $\tau$ times for the minimal exponentiation.
One drawback of the Pietrzak VDF is that as the time delay parameter increases, the proof size and the number of halving protocol repetitions for the verifier also increase. In the original paper proposing the Pietrzak VDF, a method to enhance efficiency was suggested. Since the halving protocol also consumes resources, the verifier can repeat the halving protocol $\tau-\delta$ times instead of $\tau$ times. Then, the verifier needs to check $x^{2^{2^{\delta}}}\stackrel{{\scriptstyle?}}{{=}}y$. Furthermore, this approach helps to reduce the size of the proof. Algorithm 2 describes this refined verification procedure.
In most systems, the costs of each computation and the overhead due to proof size are dependent on the specific system environment, making it difficult to determine the optimal value for the proof shortening parameter, $\delta$. However, in the EVM, the costs for data (calldata) and arithmetic computations are predefined. This led us to hypothesize that it is possible to identify the optimal $\delta$ value for implementation, thereby minimizing both computation and calldata costs in blockchain environments.
Algorithm 1 Pietrzak VDF Verification
1:input: $x,y,T,\{\pi_{i}\}_{i=1}^{\tau},N$
2:output: True or False
@@ 3.1 Pietrzak VDF verification
4:$(x_{1},y_{1})\leftarrow(x,y)$
5:for $i\leftarrow 1$ to $\tau$ do
6:     $v_{i}\leftarrow\pi_{i}$
7:     $r_{i}\leftarrow keccak256(x_{i}||y_{i}||v_{i})$
8:     $x_{i+1}\leftarrow x_{i}^{r_{i}}\cdot v_{i}\mod N$
9:     if $T/2^{i-1}$ is odd then
10:         $y_{i}\leftarrow y_{i}^{2}\mod N$
11:     end if
12:     $y_{i+1}\leftarrow v_{i}^{r_{i}}\cdot y_{i}\mod N$
13:end for
14:if $y_{\tau+1}=x_{\tau+1}^{2}$ then
15:     return True
16:else
17:     return False
18:end if
Algorithm 2 Refined Pietrzak VDF Verification
@@
3.2 Target Bit Length and Target Delay
The security of Pietrzak VDF configuration depends on the difficulty of the prime factoring problem, akin to that in the RSA encryption system. Currently, a 2048-bit key length is deemed sufficient for maintaining security standards. However, considering potential advancements in computational capabilities, a 3072-bit key length might be necessary after ten years. Therefore, our experiments included tests with both 2048-bit and 3072-bit settings for VDFs.
In a recent study on VDF implementation using C++ [7], VDF evaluation times with the time delay parameter $T$ ranging from $2^{20}$ to $2^{25}$ can be completed within one minute with a personal CPU. The evaluation involves performing $y=x^{2^{T}}$. In other words, the target range for $\tau$ is 20 – 25 where $T=2^{\tau}$. We chose to conduct our VDF experiments within the same parameter range, as it strikes a balance between computational feasibility and practical utility while avoiding excessive resource consumption.